OpenStudy (anonymous):

The lengths of the sides of a rectangle become 4 times greater than the original measurements. Which statement is true about the rectangle? Answer Its area becomes 4 times greater, and perimeter becomes 16 times greater. Its area becomes 16 times greater, and perimeter becomes 4 times greater. Its area becomes 20 times greater, and perimeter becomes 4 times greater. Its area becomes 4 times greater, and perimeter becomes 20 times greater.

OpenStudy (asnaseer):

imagine the rectangle has a length 'l' and width 'w' - what would its area be in this case?

OpenStudy (anonymous):

original area=l*w new area=4l*4w can you see it now?

OpenStudy (anonymous):

I think so, so the answer is b?

OpenStudy (anonymous):

yes, 4l*4w=16lw

OpenStudy (anonymous):

ok thank you! I see, i get it now!